OpenStudy (anonymous):

27-3+10+x = 8, what is x?

OpenStudy (lgbasallote):

just do your basic arithmetic...transpose...then do your basic arithmetic again...

OpenStudy (anonymous):

whats the answer?

OpenStudy (lgbasallote):

x = -26

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

3+10 =13. 27-13 = 14. 8-14 = oooooh I see

OpenStudy (anonymous):

thanks